OpenStudy (anonymous):

2x-8y=-24,4x+y=-2,x-4y=4 is that a perpenaicuar or not perpenaicuar

OpenStudy (aaliyahfairgood):

are you saying perpendicular

OpenStudy (aaliyahfairgood):

(2x-8y=24) -1/8 (4x+y=-2) (x-4y=4)-1 -1/4x+y=-3 -4x -4x -x+4y=-4 y=1/4x-3 y=-4x-2 (4y=x-4)/4 y=1/4x-1 y=1/4x-3 is perpendicular to y=-4x-2 y=1/4x-1is perpendicular to y=-4x-2 y=1/4x-3 is not perpendicular to y=1/4x-1